A car 2000kg tied to the rope of 15m whose end is located at the center. If the car is moving at the rate of 8 m/s. Calculate the centripetal force acting on it?

1
Expert's answer
2022-04-06T18:36:47-0400

The centripetal force

"F=\frac{mv^2}{R}=\frac{2000*8^2}{15}=8533\:\rm N"